Prove that any uniformly continuous function from a dense subset of a metric space X to a complete metric space Y has a uniformly continuous function from X to Y


1
Expert's answer
2021-07-27T10:56:27-0400

We have: (X,d) and (Y,e) are metric spaces, "f:S\to Y"

For each "\epsilon\isin R^+" , let "\delta_{\epsilon}" be the largest real number in (0,1] such that for all "a,b\isin S" (dense subset of a metric space X) with d(a, b)<"\delta_{\epsilon}" we have e(f(a), f(b))<"\epsilon" . Then, for each "x\isin X" , let "B_{x,\epsilon}=b_X[x;\delta_{\epsilon})" . Suppose "z\isin X" . Because "d(a,b)<\delta_{\epsilon}" for all "a,b\isin S\land B_{z,\epsilon}" , we have also "e(f(a),f(b))<\epsilon" , whence "diam(f(B_{z,\epsilon}))\le\epsilon" . Because S is dense in X, each of the balls "B_{z,\epsilon}" has non-empty intersection with S, so each of the sets "f(B_{z,\epsilon})" is non-empty. Because Y is complete, the nest "\{\overline{f(B_{z,\epsilon})}|\epsilon\isin R^+\}" of non-empty closed subsets of Y has singleton intersection. Set "\tilde{f}(z)" to be the sole member of that intersection; it is equal to "f(z)" if "z\isin S" , so, when this action performed for each "z\isin X" , "\tilde{f}" is an extension of f to X.

We must show that "\tilde{f}" is uniformly continuous.

Let "\gamma\isin R^+" and suppose that "u,v\isin X" satisfy "d(u,v)<\delta_{\gamma/3}/2" . Then the open set "B_{u,\gamma/3}\land B_{v,\gamma/3}" contains both u and v and is thus not empty; because S is dense in X, it contains some point a of S. Therefore "f(a)\isin f(B_{u,\gamma/3})\land f(B_{v,\gamma/3})" . By definition, we have both "\tilde{f}(u)\isin f(B_{u,\gamma/3})" and "\tilde{f}(v)\isin f(B_{v,\gamma/3})" . We have shown above that these sets have diameter not exceeding "\gamma/3" , forcing both "e(f(a),\tilde{f}(u))\le\gamma/3" and "e(f(a),\tilde{f}(v))\le\gamma/3" and yielding "e(\tilde{f}(u),\tilde{f}(v))\le 2\gamma/3<\gamma" . So "\tilde{f}" is uniformly continuous on X.